The hydrogen gas formed in a chemical reaction is collected over water at 30.0 ℃ at a total pressure of 732 mm Hg.

a) What is the partial pressure of the hydrogen gas collected in this way?

b) If the total volume of gas collected is 720 ml, what mass of hydrogen gas is collected?
Answer #1

(a) The pressure of water at 30C is 31.82mmHg.
Pressure of hydrogen=732mmHg -31.82mmHg= 700mmHg

(b) P=(700mmHg)(1atm/760mmHg)=0.921atm, V=720mL=0.72L, T=30C +273= 303K

PV=nRT, so n=PV/RT=(0.921atm)(0.72L)/(0.08206L-atm/K-mol)(303K)=0.0267mol

Molecular weight of H2: 2.02g/mol

(0.0267mol)(2.02g/mol)=0.054g